Two wires A and B are of the same material. Their lengths are in the ratio 1:2 and the diameter are in the ratio 2 : 1. If they are pulled by the same force, then increase in length will be in the ratio


        (a)    2:1


        (b)    1:4


        (c)    1:8


        (d)    8:1


Grade:upto college level

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
elongation is proportional to lenght / Area
=>lenght/radius^2
=>elong. A/elong. B=1*4/2*1=2/1=2:1
Thanks & Regards
Arun Kumar, Btech, IIT delhi
Askiitians Faculty